# UDEEP: Edge-based Computer Vision for In-Situ Underwater Crayfish and Plastic Detection

### v1

@@ 5 Conclusions and Future Work
This article presents the Underwater DL edge computing platform (UDEEP), a novel monitoring platform for underwater crayfish and plastics identification that is world-leading in terms of originality, significance and rigour. The platform integrates diverse methods, including background subtraction for selectively cropping regions of interest in high-resolution images to detect crayfish and underwater plastic debris. By providing accurate data on the presence, distribution, and population density of these species, the UDEEP platform makes a substantial contribution to monitoring initiatives and aids in addressing the spread of invasive species A public custom underwater crayfish dataset, consisting of 2486 images Machado et al. (2023) was generated as part of this Research Project and made publicly available. Additionally, a second public custom dataset, also created by the authors, focuses on plastic and comprises 1220 images Machado (2022) was generated and released into the public domain.
This paper further presents experiments using YOLO v5 nano and small and YOLO v8 nano and small image classification algorithms that were trained to identify crayfish and plastic in underwater environments using a NJN board. From the results, it is evident that the two versions of YOLO were able to detect and count the number of crayfish and plastic with a high degree of accuracy. The results of measuring the inference current showed that YOLO v8s required the least amount of current and YOLO v5n required the most amount of current among the four models.
Future work involves integrating the models into other versions of edge computing devices for example the NVIDIA TX2 which will significantly enhance the generation of statistics related to the detection, quantity, and location of invasive and endangered species. The UDEEP platform will empower policymakers, conservation authorities, and researchers with real-time data on the spread of invasive non-native species and discarded plastic debris floating along water bodies. The availability of such data will facilitate the enactment of well-informed conservation laws and the implementation of effective mitigation measures to safeguard native endangered species and reduce plastic pollution.
@@
While the current study concentrated on identifying discarded plastic and signal crayfish as the non-native species responsible for the decimation of native white-clawed crayfish, the UDEEP platform has the potential to be further trained to recognise other non-native species. For instance, it could be extended to identify invasive non-native species such as the red-eared terrapin and Chinese mitten crab, broadening the scope of its applicability and environmental impact assessment.

### v2

@@ 6 Conclusions and Future Work
The Underwater DL edge computing platform (CED), a novel monitoring platform for underwater crayfish and plastics identification that is world-leading in terms of originality, significance and rigour is presented in this article. The platform integrates diverse methods, including background subtraction for selectively cropping regions of interest in high-resolution images to detect crayfish and underwater plastic debris. By providing accurate data on the presence, distribution, and population density of these species, the CED platform makes a substantial contribution to monitoring initiatives and aids in addressing the spread of invasive species A public custom underwater crayfish dataset, consisting of 2486 images [machado_2023] was generated as part of this Research Project and made publicly available. Additionally, a second public custom dataset, also created by the authors, focuses on plastic and comprises 1220 images [machado_2022] was generated and released into the public domain.
YOLO v5s demonstrated the highest overall detection accuracy (mAP@0.5 of 0.90) and precision (0.93), demonstrating its strong capability to identify crayfish and plastic debris with minimal false positives. YOLO v8s achieved the best recall of 0.84 and localisation precision, achieving the lowest validation box loss of 0.95. Although YOLO v5n and YOLO v8n also performed well with mAP@0.5 scores of 0.87 and 0.86 respectively, the YOLO v5 variants generally showed higher residual errors and notably slower convergence, with YOLO v5n being the slowest to stabilise. For applications on resource-constrained embedded systems like the NJON, YOLO v8n emerged as the most optimal model when considering a balance between performance and operational efficiency. On the GPU, it recorded the fastest overall inference time at 0.29 seconds and the lowest total energy consumption at just 0.09 NJON, significantly outperforming alternatives like YOLO v5n (0.4829 seconds and 0.40 J). Furthermore, YOLO v8n consistently exhibited exceptionally low Mean SOC power across various test conditions, reaching as low as 0.0001 W under Gaussian Noise, further underscoring its hardware-level power optimisation. CPU inference times were notably slower across all models (e.g., YOLO v8n at 0.59 seconds vs. YOLO v8s at 1.31 seconds), reinforcing the necessity of GPU acceleration for real-time applications. In summary, while YOLO v5s excelled in overall detection accuracy and precision, and YOLO v8s offered strong recall and precise localisation, YOLO v8n is identified as the most suitable model for real-time embedded applications. Its efficiency, highlighted by its leading GPU inference speed and lowest energy consumption, combined with its robust performance and excellent SOC power management, directly addresses the demanding operational requirements of underwater autonomous systems.
Future work will focus on equipping CED with pose estimation capabilities to enable detailed analysis of fish behaviour and movement patterns, which can serve as indicators of health and stress. Abnormal swimming patterns, changes in fin position, or irregular body curvature detected through pose estimation may signal injuries, disease, or environmental distress. When combined with contextual sensor data such as water quality, temperature, and oxygen levels, this integrated approach will enhance CED’s ability to monitor aquatic animal health and behaviour in real time, supporting conservation, research, and early intervention strategies. The CED platform will empower policymakers, conservation authorities, and researchers with real-time data on the spread of invasive non-native species and discarded plastic debris floating along water bodies. The availability of such data will facilitate the enactment of well-informed conservation laws and the implementation of effective mitigation measures to safeguard native endangered species and reduce plastic pollution.
@@
While the current study concentrated on identifying discarded plastic and signal crayfish as the non-native species responsible for the decimation of native white-clawed crayfish, the CED platform has the potential to be further trained to recognise other non-native species. For instance, it could be extended to identify invasive non-native species such as the red-eared terrapin and Chinese mitten crab, broadening the scope of its applicability and environmental impact assessment.
Author Contributions
D.M. led the implementation of the Cognitive Edge Device (CED) platform, developed the machine learning algorithms, performed the experiments, and conducted the results analysis. F.F.T. contributed to system development, supported the implementation of the detection pipeline, and assisted in analysing and interpreting the experimental outcomes. J.J.B. provided methodological guidance, contributed to refining the machine learning workflow, and supported the review of experimental results. A.L. contributed to conceptual oversight, provided feedback on system design and methodology, and participated in reviewing and improving the manuscript. S.W.Y. assisted with data preparation, supported the experimental setup, and contributed to reviewing the manuscript for technical accuracy. M.M.H. contributed to the interpretation of results, supported the preparation of figures and tables, and provided critical revisions to the manuscript. P.S. contributed to hardware integration, including sensor and embedded platform configuration, and supported the evaluation of system performance in edge deployment scenarios. P.M. conceived and supervised the study, guided the development of the CED platform, contributed to the research design, assisted in results analysis, and led the writing and revision of the manuscript.
All authors reviewed and approved the final manuscript.
Funding statement
This research was supported by QR funding from Nottingham Trent University’s Department of Computer Science
Conflict of Interest
The authors declare that there are no known competing financial interests or personal relationships that could have appeared to influence the work reported in this article.
Data Availability
The source code developed for the CED platform, including scripts for dataset preprocessing, model training, inference, and edge deployment, is openly available at https://github.com/denomon/CognitiveEdgeDeviceForRTEMonitoring. The underwater datasets used in this study are publicly accessible. The crayfish dataset comprising 2,486 annotated underwater images is available on Zenodo at [machado_2023] and the underwater plastic debris dataset, consisting of 1,220 annotated images, is available on Zenodo at [machado_2022]. All datasets and source code used in this work are freely accessible for replication, benchmarking, and further research.
Clinical trial number
Not applicable.
Consent to Publish declaration
Not applicable.
Ethics and Consent to Participate declarations
Not applicable.